OpenStudy (anonymous):

The equation of the line that goes through the point (3,4) and is parallel to the line going through the points (-2,3) and (3,2) can be written in the form y=mx+b where m=?? and b=??

OpenStudy (saifoo.khan):

parallel means same slope.

OpenStudy (anonymous):

find the slope of the line that goes through (-2,3) and (3,2) by the formula m=(y1-y2)/(x1-x2), and then insert that slope into y=mx+b. to find b, plug your coordinates (3,4) into y and x to figure out the y-intercept.

OpenStudy (andrew314):

slope (m) = -1/5 intercept (b) = 23/5

OpenStudy (anonymous):

thanks for clarifying!!